GoatyGoat:

What is 18m^4 - 10m^2?

curriful:

how do u want it to be solved? u want it factored or smth?

GoatyGoat:

@curriful wrote:
how do u want it to be solved? u want it factored or smth?
tbh idk my math teacher just told me what is 18m^4 - 10m^2 ._.

toga:

18m^4 - 10m^2 can be simplified using the common factor of 2m^2, which yields: 18m^4 - 10m^2 = 2m^2(9m^2 - 5)

GoatyGoat:

@toga wrote:
18m^4 - 10m^2 can be simplified using the common factor of 2m^2, which yields: 18m^4 - 10m^2 = 2m^2(9m^2 - 5)
ok ty

toga:

@goatygoat wrote:
@toga wrote:
18m^4 - 10m^2 can be simplified using the common factor of 2m^2, which yields: 18m^4 - 10m^2 = 2m^2(9m^2 - 5)
ok ty
np